OpenStudy (anonymous):

what two numbers does square root 117 lie

OpenStudy (anonymous):

100 < 117 < 121 Taking square root \[\sqrt{100} < \sqrt{117} < \sqrt{121}\] 10 < sqr. root 117 < 11 Square root of 117 lies between 10 and 11.

OpenStudy (anonymous):

ok

OpenStudy (goformit100):

Warm Welcome To Open Study ......... Sir/ Ma'am @msvanny